OpenStudy (faithanna101):

A 7 1/4 in. circular power saw rotates at 5200 revolutions a min. Find the angular speed; radians a min.

OpenStudy (championcult):

\[w=\frac{ \theta }{ t }\] => \[\frac{ 5200 rad }{ 1 \min } * \frac{ 2\Pi }{ 1 rev } * \frac{ 1 \min }{ 60 \sec }\] => \[\frac{ 10400\Pi rad }{ 60 \sec }\] => \[\frac{ 520\Pi }{ 2 } rad per \sec\] => => => V=rw => \[7\frac{ 1 }{ 4 } (\frac{ 520\Pi }{ 2 }) \] => \[7.25(260\Pi) \] => \[1885\Pi \] => \[5922 cm per \sec\]